OpenStudy (anonymous):

is x=6 extranous or no

OpenStudy (anonymous):

@taramgrant0543664

OpenStudy (xapproachesinfinity):

where is the equation?

OpenStudy (xapproachesinfinity):

you want me to make one lol

OpenStudy (xapproachesinfinity):

ok good!! what did you do thus far?

OpenStudy (anonymous):

igt x=6

OpenStudy (anonymous):

but ineed to know if its extranous or not

OpenStudy (xapproachesinfinity):

well put it back into the equation and see if it does work? if we put six what is the problem?

OpenStudy (xapproachesinfinity):

It does not really work does it?

OpenStudy (xapproachesinfinity):

that will produce division by zero yes? which is not allowed

OpenStudy (xapproachesinfinity):

you agree or disagree? what is your stance?

OpenStudy (xapproachesinfinity):

hhhh well can we divide by 0 can we have 6/0 in the left in the right 6/0-6/2

OpenStudy (xapproachesinfinity):

what is 6/0? can you have an answer?

OpenStudy (anonymous):

not so not extanrous

OpenStudy (xapproachesinfinity):

ok once step at a time what is extraneous solution

OpenStudy (xapproachesinfinity):

let me answer this, it is a solution that you find by correct algebraic maneuvers but it does not work for the equation when we put it back

OpenStudy (anonymous):

An extraneous solution is a root of a transformed equation that is not a root of the original equation because it was excluded from the domain of the original equation.

OpenStudy (xapproachesinfinity):

yeah that's the definition right there so from the start x=6 is excluded from domain yes

OpenStudy (xapproachesinfinity):

but when you were solving the equation to transform it using some maneuvers then you got x=6

OpenStudy (xapproachesinfinity):

that's why we say it is extraneous solution because it is not included in the domain

OpenStudy (xapproachesinfinity):

here is what we do \[\frac{x}{x-6}=\frac{x}{x-6}-\frac{6}{2}\] domain here is every real number except 6

OpenStudy (xapproachesinfinity):

slight mistake first should be 6/x-6

OpenStudy (xapproachesinfinity):

so \[\frac{6}{x-6}=\frac{x}{x-6}-\frac{6}{2}\] \[\frac{6}{x-6}=\frac{2x-6}{2(x-6)}\] \[\frac{6}{\cancel {x-6}}=\frac{2x-6}{2\cancel{(x-6)}}\] \[6=\frac{2x-6}{2}\Longrightarrow 12 =2x-6 \Longrightarrow x=9\]

OpenStudy (xapproachesinfinity):

how did you got x=6?

OpenStudy (xapproachesinfinity):

i got 9

OpenStudy (xapproachesinfinity):

oh my bad i miss something

OpenStudy (xapproachesinfinity):

correct the solution found is x=6 just missed up above in line 2

OpenStudy (xapproachesinfinity):

so 6 we said excluded therefore it is extraneous

OpenStudy (xapproachesinfinity):

was that helpful? clear?